# MeshAnything V2: Artist-Created Mesh Generation With Adjacent Mesh Tokenization

### v1

@@ 3.2 MeshAnything V2
Following (Chen et al., 2024b), MeshAnything V2 also targets generating AMs aligned to a given shape, allowing integration with various 3D asset production pipelines to achieve highly controllable AM generation. That is, we aim to learn the distribution: $p(\mathcal{M}|\mathcal{S})$, where $\mathcal{M}$ represents the AM and $\mathcal{S}$ represents the 3D shape condition.
As in (Chen et al., 2024b), V2 uses point clouds as the shape condition input $\mathcal{S}$. We also use the same point cloud–Artist-Created Mesh data pairs $(\mathcal{M},\mathcal{S})$ collected in (Chen et al., 2024b). The target distribution $p(\mathcal{M}|\mathcal{S})$ is learned with a decoder-only transformer with the same size and architecture as in (Chen et al., 2024b). To inject $\mathcal{S}$ into the transformer, we first encode it with a pretrained point cloud encoder (Zhao et al., 2024) into a fixed-length token sequence $\mathcal{T}_{S}$ and then set it as the prefix of the transformer’s token sequence. We then process paired $\mathcal{M}$ into mesh token sequence $\mathcal{T}_{M}$. It is concatenated to the point cloud token sequence as the transformer’s ground truth sequence. After training the transformer with cross-entropy loss, we input $\mathcal{T}_{S}$ and let the transformer autoregressively generate the corresponding $\mathcal{T}_{M}$, which is then detokenized into $\mathcal{M}$.
The key difference between (Chen et al., 2024b) and our method is the way we obtain $\mathcal{T}_{M}$. Instead of the naive mesh tokenization method used in (Chen et al., 2024b), we process $\mathcal{M}$ with the newly proposed Adjacent Mesh Tokenization (AMT) and obtain a more compact and efficient sequence $\text{Seq}_{V}$. Following (Chen et al., 2024a), we discard the VQ-VAE and directly use the discretized coordinates from $\text{Seq}_{V}$ as token indices. We then add a newly initialized codebook entry to represent the & in the AMT sequence. Finally, we sequentially combine the coordinate token sequence and the special token for & to obtain the mesh token sequence $\mathcal{T}_{M}$ for transformer input. As mentioned in Sec. 3.1, it is worth noting that whether or not VQ-VAE is used does not affect the application and effectiveness of AMT.
@@
V2 W.O. AMT 1.454 5.867 0.913 297.4 561.6 1.132 1.124
V2 0.802 4.587 0.935 314.7 590.7 1.207 1.192
To facilitate the transformer’s learning of the sequence patterns of AMT, in addition to the absolute positional encoding used in (Chen et al., 2024b), we add the following embeddings for Adjacent Mesh Tokenization (AMT): when representing a face with three vertices, we add a specific embedding for the three new vertices; when representing a face with one vertex, we add a different embedding for the single new vertex. Additionally, we provide a distinct embedding for the & token.
@@ 4.3 Quantitative Experiments
Averaged Token Sequence Length. We randomly sample $10k$ mesh samples with fewer than $1600$ faces from our dataset and tokenize them into token sequences using the previous tokenization method and AMT separately. In transformer Vaswani et al. (2017) learning, a vertex requires several tokens to represent it, whereas the special symbol & only requires one token. Therefore, the loss incurred by AMT when it cannot find the next adjacent face is relatively small. Since three tokens are used to represent a vertex in (Chen et al., 2024b; a) as well as in this paper, we calculate the final input token length for the transformer based on this standard. For each sample, we divide the length of the token sequence obtained by our method by the length obtained by the previous method, and then average the results to determine the average reduction in token sequence length achieved by AMT. Across the aforementioned $10k$ mesh samples, we found this ratio to be $0.4973$, indicating that the AMT method significantly reduces length of token sequence. Additionally, this experiment only demonstrates the superiority of the AMT algorithm in shortening sequence length; its more compact and well-structured sequence characteristics also provide further advantages for sequence learning.
Ablation Study. We ablate the effectiveness of AMT by comparing the results of MeshAnything V2 with its variant without AMT. The variant follows exactly the same settings as V2, except that AMT is replaced with naive mesh tokenization as in previous methods (Siddiqui et al., 2023; Chen et al., 2024a; b). This also serves as a fairer comparison with (Chen et al., 2024b), as the original model of (Chen et al., 2024b) was trained on data with fewer than $800$ faces rather than $1600$. Moreover, (Chen et al., 2024b) is trained on a small total batch size of $64$ while we observed a noticeable performance improvement when the batch size was increased to V2’s $256$.
We follow the evaluation metric settings of (Chen et al., 2024b) to quantitatively assess mesh quality. We uniformly sample $100K$ points from both the ground truth meshes and the generated meshes and compute the following metrics on our $2$K evaluation dataset:
@@
As shown in Table 1, our results indicate that V2 significantly outperforms its variant, demonstrating the effectiveness of AMT. It shows that AMT not only improves training speed and reduces memory pressure but also enhances generation quality. Notably, although V2 and its variant were trained for the same number of iterations, the variant consumed nearly two times the GPU hours of V2. Additionally, the table shows that AMT takes slightly more vertices and faces. A likely reason for this is that the variant’s performance is still too weak, and it tends to ignore details and use a simpler topology when representing complex meshes with high face counts. We also observe that both AMT and the variant have vertex and face ratios greater than $1.0$, meaning they use more faces on average relative to the ground truth, unlike the results in (Chen et al., 2024b), which were less than $1$ (around $0.88$). We suspect this is because the dataset for V2 contains more complex meshes with over $800$ faces, which causes the model to occasionally produce more complex topologies for simple shapes.
5 Limitations and Conclusion
@@
In this work, we present MeshAnything V2, a shape-conditioned Artist-Created Mesh (AM) generation model that generates AM aligned to given shapes. V2 significantly outperfroms MeshAnything (Chen et al., 2024b) in both performance and efficiency with our newly proposed Adjacent Mesh Tokenization (AMT). Different from previous methods that use three vertices to represent a face, AMT uses a single vertex whenever possible. Our experiments demonstrate that AMT averagely reduces the token sequence length by half. The compact, and well-structured token sequence from AMT greatly enhances sequence learning, thereby significantly improving the efficiency and performance of AM generation.
@@
Limitations. Although there is a large improvement over V1, the accuracy of MeshAnything V2 is still insufficient for industrial applications. More efforts are needed to improve the model’s stability and accuracy.

### v2

@@ 3.2 MeshAnything V2
Following [5], MeshAnything V2 also targets generating AMs aligned to a given shape, allowing integration with various 3D asset production pipelines to achieve highly controllable AM generation. That is, we aim to learn the distribution: $p(\mathcal{M}|\mathcal{S})$, where $\mathcal{M}$ represents the AM and $\mathcal{S}$ represents the 3D shape condition.
As in [5], V2 uses point clouds as the shape condition input $\mathcal{S}$. We also use the same point cloud–Artist-Created Mesh data pairs $(\mathcal{M},\mathcal{S})$ collected in [5]. The target distribution $p(\mathcal{M}|\mathcal{S})$ is learned with a decoder-only transformer with the same size and architecture as in [5]. To inject $\mathcal{S}$ into the transformer, we first encode it with a pretrained point cloud encoder [33] into a fixed-length token sequence $\mathcal{T}_{S}$ and then set it as the prefix of the transformer’s token sequence. We then process paired $\mathcal{M}$ into mesh token sequence $\mathcal{T}_{M}$. It is concatenated to the point cloud token sequence as the transformer’s ground truth sequence. After training the transformer with cross-entropy loss, we input $\mathcal{T}_{S}$ and let the transformer autoregressively generate the corresponding $\mathcal{T}_{M}$, which is then detokenized into $\mathcal{M}$.
The key difference between [5] and our method is the way we obtain $\mathcal{T}_{M}$. Instead of the naive mesh tokenization method used in [5], we process $\mathcal{M}$ with the newly proposed Adjacent Mesh Tokenization (AMT) and obtain a more compact and efficient sequence $\text{Seq}_{V}$. Following [4], we discard the VQ-VAE and directly use the discretized coordinates from $\text{Seq}_{V}$ as token indices. We then add a newly initialized codebook entry to represent the & in the AMT sequence. Finally, we sequentially combine the coordinate token sequence and the special token for & to obtain the mesh token sequence $\mathcal{T}_{M}$ for transformer input. As mentioned in Sec. 3.1, it is worth noting that whether or not VQ-VAE is used does not affect the application and effectiveness of AMT.
@@
V2 W/O AMT 0.895 4.832 0.924 302.4 556.7 1.105 1.062 1.000
V2 0.874 4.721 0.933 308.6 571.8 1.127 1.097 0.497
To facilitate the transformer’s learning of the sequence patterns of AMT, in addition to the absolute positional encoding used in [5], we add the following embeddings for Adjacent Mesh Tokenization (AMT): when representing a face with three vertices, we add a specific embedding for the three new vertices; when representing a face with one vertex, we add a different embedding for the single new vertex. Additionally, we provide a distinct embedding for the & token.
Face Count Condition. Considering that some applications require control over the approximate face count, we explored the addition of a face count condition in mesh generation. We initialized an embedding book with a size equal to the maximum allowed number of faces. Based on the current face count in the mesh, we retrieve the corresponding embedding from this book and place it after the point cloud prefix to represent the face count. During training, we add a random variable to the face count to introduce some variation and prevent overfitting to an exact condition. During training, we add a random variable to the face count to introduce some variation and prevent overfitting to an exact condition. Additionally, we drop this condition with a 10% probability to further improve condition robustness.
Masking Invalid Predictions. PolyGen [15] introduces Masking Invalid Predictions during inference. In PolyGen, vertices are generated first, followed by faces, and both vertices and faces are sorted. Without applying constraints during inference, generative models may produce results that do not follow the sequence structure, such as generating faces before completing all vertices or generating vertices that violate the sorted coordinates order. PolyGen [15] addresses this issue by masking invalid logits during the inference phase, ensuring that only valid results are generated.
We followed this design in our PolyGen experiments. Due to the widespread use of coordinate sorting in meshes, this design can also be adopted in other tokenization methods. In the AMT experiments, we incorporated this design as well. For example, we enforced a rule in AMT where, when starting a new strip, at least three vertices must be generated before allowing any interruptions.
@@ 4.1 Implementation Details
For dataset preparation, we adopt a similar dataset preparation technique as used in [5], with several modifications. Firstly, we noticed that Objaverse contains some duplicate mesh data, where the meshes are identical in shape but differ in texture. We detected these duplicates by calculating the differences in vertex coordinates and retained only one instance of each mesh. Additionally, in [5], the ground truth meshes were first processed into SDF using mesh2sdf [24]. While mesh2sdf occasionally produces some failure cases, we filtered out these failures using Chamfer Distance. Different from the dataset used in [5], we used meshes with fewer than $1600$ faces as the experimental dataset, compared to $800$ in the previous work, and also included a small portion of data from ObjaverseXL. This resulted in a dataset containing $230$K point clouds and mesh pairs. We randomly sampled $4$K data samples as the evaluation dataset.
To accommodate meshes with more faces, we sampled $8192$ points instead of $4096$ for each point cloud. Besides, unlike [5], we update the point encoder from [33] during training because we find its accuracy insufficient for handling complex meshes with up to $1600$ faces.
@@ 4.2 Qualitative Experiments
Table 2: Comparison of Tokenization Methods. We compare various mesh tokenization methods and their impact on the generated mesh results in mesh generation. Note that the metrics are calculated by sampling 10K points cloud on each mesh.
[…]
(×$\times 10^{-2}$) (×$\times 10^{-2}$)
Baseline 2.478 18.21 0.893 95.4 178.9 0.974 0.940 1.000 1.150
Unsort 8.151 31.86 0.794 117.4 213.1 1.189 1.219 1.000 1.234
PolyGen(AMT) 3.226 22.97 0.872 93.2 172.2 0.928 0.936 0.372 1.589
AMT 2.348 19.33 0.904 102.7 187.1 1.013 0.983 0.492 1.363
AMT(Swap) 2.517 19.86 0.913 110.8 198.9 1.098 1.102 0.455 1.416
@@ 4.3 Quantitative Experiments
Evaluation Metrics. We follow the evaluation metric settings of [5] to quantitatively assess mesh quality. We uniformly sample point clouds from both the ground truth meshes and the generated meshes and compute the following metrics on our $2$K evaluation dataset:
@@
  • Perplexity: We report training perplexity to measure how well the mesh generation model predicts a sample. A lower perplexity indicates more accurate predictions.
  • Sequence Ratio (S_Ratio): The sequence length compression ratio of the used mesh tokenization methods.
Ablation Study. We ablate the effectiveness of AMT by comparing the results of MeshAnything V2 with its variant without AMT. The variant follows exactly the same settings as V2, except that AMT is replaced with naive mesh tokenization as in previous methods [19, 4, 5]. This also serves as a fairer comparison with [5], as the original model of [5] was trained on data with fewer than $800$ faces rather than $1600$. Moreover, [5] is trained on a small total batch size of $64$ while we observed a noticeable performance improvement when the batch size was increased to V2’s $256$. We sample 100K points on each mesh to calculate the aforementioned metrics.
@@
As shown in Table 1, our results indicate that V2 significantly outperforms its variant, demonstrating the effectiveness of AMT. It shows that AMT not only improves training speed and reduces memory pressure but also enhances generation quality. Notably, although V2 and its variant were trained for the same number of iterations, the variant consumed nearly two times the GPU hours of V2. Additionally, the table shows that AMT takes slightly more vertices and faces. A likely reason for this is that the variant’s performance is still too weak, and it tends to ignore details and use a simpler topology when representing complex meshes with high face counts. We also observe that both AMT and the variant have vertex and face ratios greater than $1.0$, meaning they use more faces on average relative to the ground truth, unlike the results in [5], which were less than $1$ (around $0.88$). We suspect this is because the dataset for V2 contains more complex meshes with over $800$ faces, which causes the model to occasionally produce more complex topology for simple shapes.
Comparison of Tokenization Methods. Below, we list the methods to be compared in the following sections, along with their specific implementation details. In the tokenization methods comparison experiments, we use the smaller OPT-125M [32] and select meshes with up to 400 faces from the aforementioned datasets as the experimental dataset. We sample 10K points on each mesh to calculate the aforementioned metrics.
  • Baseline. This refers to the naive mesh tokenization method used in previous works, where each face is strictly represented by three vertices.
  • PolyGen(AMT). This is a combination of the tokenization method from PolyGen [15] with AMT. Polygen first generates vertex coordinates and then expresses the faces by generating vertex indices. We discuss this setting in detail in the Supplementary Materials.
  • AMT. This refers to the Adjacent Mesh Tokenization proposed in Sec. 3.1, as shown in Figure Aug1.
  • AMT(Swap). As described in Sec. 3.1, this method combines the AMT algorithm with vertices swap. An additional special token is used to represent the swap operation.
  • Unsort. This is based on the naive mesh tokenization method from previous works [15, 19], but without sorting the mesh. Instead, the mesh is used in the order in which it is stored. It is important to note that the meshes in Objaverse [6] are created by human, and their storage may already include some level of inherent order. The purpose of this method is to compare with the Baseline and verify the importance of sorting.
The results are shown in Tab. 2. Unsort vs Baseline comparison validates the importance of tokenization method regularity for mesh generation.
PolyGen(AMT) [15] achieves the best S_Ratio (sequence length compression ratio), indicating its high efficiency. However, there is a noticeable performance gap compared to Baseline and AMT, indicating the produced mesh quality is not high enough. The primary reason for this might be that the autoregressive model struggles to interact accurately with previously generated vertices when producing faces, which increases the difficulty of sequence learning.
The results from AMT demonstrates that it shortens the token length without compromising mesh quality, as confirmed by its comparison with the Baseline. Additionally, in AMT(Swap), the introduction of the Swap special token further increases the compression ratio without affecting mesh quality.
From the above experiments, we can conclude that both the regularity of the token sequence and the compression ratio are crucial. Although PolyGen [15] achieves a better compression ratio, its token sequence organization is not well-suited for sequence learning, making it a sub-optimal mesh tokenization method compared to AMT. In contrast, both AMT and AMT(Swap) reduce the compression ratio without increasing the difficulty of sequence learning, making them more suitable for mesh generation.
5 Conclusion
@@
In this work, we present MeshAnything V2, a shape-conditioned Artist-Created Mesh (AM) generation model that generates AM aligned to given shapes. V2 significantly outperfroms MeshAnything [5] in both performance and efficiency with our newly proposed Adjacent Mesh Tokenization (AMT). Different from previous methods that use three vertices to represent a face, AMT uses a single vertex whenever possible. Our experiments demonstrate that AMT averagely reduces the token sequence length by half. The compact, and well-structured token sequence from AMT greatly enhances sequence learning, thereby significantly improving the efficiency and performance of AM generation.
Appendix A Additional Experiments
We provide additional quantitative experiments and detailed discussion about Polygen [15] tokenization method here.
A.1 Experiments on Face Count Condition.
We tested the effectiveness of the face count condition by scaling the ground truth face count value by a scalar. We used the V2 model trained on the mesh dataset with fewer than 1600 faces as the test model and sampled 10K point clouds for each mesh to calculate the metrics. During inference, for each ground truth mesh, we input the scaled face count and the paired point cloud into the model and measured the effect on face count and mesh quality. As shown in Tab. 3, at a scale ratio of 0.8, the model significantly reduced the face count while maintaining mesh quality. However, at a scale ratio of 0.6, the face count did not decrease further, indicating that while the model has some ability to follow the face condition, it prioritizes mesh quality when the condition becomes difficult to meet. Similarly, when the scale ratio is set to greater than 1, the model exhibits similar behavior.
A.2 User Study.
We conducted a user study on mesh generation quality to compare MeshAnything [5] and our works. For a fair comparison, we randomly sampled 30 meshes with fewer than 800 faces from the evaluation dataset and input their paired point clouds into each model. Users were asked to select the result they preferred from the two options. We collected responses from a total of 43 users, and the voting rates for [5] and MeshAnything V2 were 32.2% and 67.8%, respectively. This indicates that the results generated by V2 are more aligned with human preference.
A.3 Discussion on PolyGen Tokenization Method.
Polygen[15] introduces a mesh tokenization approach that differs from other existing methods [19] for mesh generation. It first employs an autoregressive vertex model to generate the 3D coordinates of the mesh’s vertices. These vertices are then used as a prefix and fed into another autoregressive face model, which connects these vertices into faces, thus constructing the entire mesh. Since the 3D coordinates are already provided by the vertex model, the face model does not need to estimate the 3D coordinates, but only specifies the connections using vertex indices, significantly reducing the sequence length. During the face generation, this tokenization method can be combined with AMT further shorten the token sequence length.
Table 3: Experiments on Face Count Condition. We control the face count condition using the scale ratio, where 1.0 indicates using the ground truth face count as the condition. The experiments show that our face count condition has the ability to control the number of faces.
[…]
1.0 1.768 1.127 1.097
0.8 1.734 0.928 0.912
0.6 1.822 0.902 0.882
1.2 1.814 1.282 1.248
1.4 1.920 1.271 1.252
By using vertex indices to represent faces, PolyGen [15] tokenization consumes only one token to define a face, whereas other methods require three tokens to represent a single vertex. Assuming a vertex is referenced $n$ times, PolyGen’s tokenization requires $3+n$ tokens, whereas other tokenization methods would require $3\times n$ tokens. Although PolyGen spends one additional token when a vertex is referenced only once, in most cases, each vertex is referenced multiple times, allowing PolyGen’s approach to save sequence length.
Although PolyGen tokenization effectively reduces the token sequence length, this generation method requires the model to accurately predict vertex positions first, which may not be ideal for an autoregressive model.
In the experiments section of our main paper, we compare Polygen tokenization method with other methods. To make a fair comparison, we merge the two stage generation process of PolyGen [15] into a single model that first generates vertex coordinates and then expresses the faces by generating vertex indices. Besides, we combine PolyGen’s face generation stage with AMT to further reduce the token sequence length.
Appendix B Limitations.
@@ Appendix A Additional Experiments
Although there is a large improvement over V1, the accuracy of MeshAnything V2 is still insufficient for industrial applications. More efforts are needed to improve the model’s stability and accuracy.